# Temporal Overlapping Prediction: A Self-supervised Pre-training Method for LiDAR Moving Object Segmentation

### v1

@@ 5.2 Few-shot Results on nuScenes
We perform few-shot MOS experiments on the nuScenes dataset. All methods are first pre-trained on full sequences from the nuScenes train split. Then the pre-trained encoder of the last epoch is chosen and fine-tuned with labeled MOS data using binary Cross-Entropy loss. We randomly select 5%/10%/20%/50% of sequences from the nuScenes train split for MOS fine-tuning. To assess the convergence performance, all models are trained for a sufficient number of epochs until convergence. As it is not possible to assign weights to the two metrics to select a single “best” model, we can only choose two best models separately based on each metric. We present the test results on full sequences from the nuScenes val split in Tab. 1.
Our pre-training method improves training-from-scratch baseline across all fine-tuning data amounts. Our method achieves significant improvement in $\text{mIoU}_{\text{obj}}$, up to 12.17% relative improvement from 24.98% to 28.03% $\text{mIoU}_{\text{obj}}$ when using 10% data, while the improvement in $\text{IoU}_{\text{w/o}}$ is less pronounced. This is because we detect more moving objects with less number of points (such as pedestrians or cycles), which account for a small proportion of the total points. It is noteworthy that UnO also shows some increases on $\text{mIoU}_{\text{obj}}$, while the $\text{IoU}_{\text{w/o}}$ all decreases. This result indicates that the performance of pre-training by UnO is not consistent, and it introduces more false negatives when detecting moving objects. This will lead to erroneous moving object alerts in MOS applications and cause incorrect path planning. In contrast, our method demonstrates consistent results in most cases. We visualize some results in Fig. 4.
Data Pretrain Best $\textbf{mIoU}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
$\textbf{mIoU}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{mIoU}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
@@
Uno $\text{25.23}^{\text{{\color[rgb]{0,0.88,0}+0.96}}}$ $\text{34.31}^{\text{{\color[rgb]{0.68,0,0}-0.65}}}$ $\text{22.62}^{\text{{\color[rgb]{0.68,0,0}-0.58}}}$ $\text{34.90}^{\text{{\color[rgb]{0.68,0,0}-0.15}}}$
Ours $\textbf{25.75}^{\textbf{{\color[rgb]{0,0.88,0}+1.48}}}$ $\textbf{35.94}^{\textbf{{\color[rgb]{0,0.88,0}+0.98}}}$ $\textbf{25.75}^{\textbf{{\color[rgb]{0,0.88,0}+2.56}}}$ $\textbf{35.94}^{\textbf{{\color[rgb]{0,0.88,0}+0.90}}}$
@@
ALSO $\text{27.32}^{\text{{\color[rgb]{0,0.88,0}+2.34}}}$ $\text{35.74}^{\text{{\color[rgb]{0.68,0,0}-0.70}}}$ $\text{22.47}^{\text{{\color[rgb]{0.68,0,0}-0.51}}}$ $\textbf{38.33}^{\textbf{{\color[rgb]{0,0.88,0}+0.27}}}$
@@
Uno $\text{26.38}^{\text{{\color[rgb]{0,0.88,0}+1.40}}}$ $\text{36.56}^{\text{{\color[rgb]{0,0.88,0}+0.12}}}$ $\text{22.00}^{\text{{\color[rgb]{0.68,0,0}-0.98}}}$ $\text{37.17}^{\text{{\color[rgb]{0.68,0,0}-0.89}}}$
Ours $\textbf{28.03}^{\textbf{{\color[rgb]{0,0.88,0}+3.04}}}$ $\textbf{36.95}^{\textbf{{\color[rgb]{0,0.88,0}+0.51}}}$ $\textbf{24.69}^{\textbf{{\color[rgb]{0,0.88,0}+1.72}}}$ $\text{38.06}^{\text{{\color[rgb]{0,0.88,0}+0.00}}}$
@@
ALSO $\text{26.68}^{\text{{\color[rgb]{0,0.88,0}+1.08}}}$ $\text{40.84}^{\text{{\color[rgb]{0.68,0,0}-3.41}}}$ $\text{25.81}^{\text{{\color[rgb]{0,0.88,0}+1.28}}}$ $\text{42.33}^{\text{{\color[rgb]{0.68,0,0}-2.72}}}$
4Docc $\text{26.01}^{\text{{\color[rgb]{0,0.88,0}+0.42}}}$ $\text{42.21}^{\text{{\color[rgb]{0.68,0,0}-2.04}}}$ $\text{24.58}^{\text{{\color[rgb]{0,0.88,0}+0.06}}}$ $\text{43.11}^{\text{{\color[rgb]{0.68,0,0}-1.94}}}$
Uno $\text{28.14}^{\text{{\color[rgb]{0,0.88,0}+2.55}}}$ $\text{43.63}^{\text{{\color[rgb]{0.68,0,0}-0.62}}}$ $\text{26.78}^{\text{{\color[rgb]{0,0.88,0}+2.26}}}$ $\text{43.97}^{\text{{\color[rgb]{0.68,0,0}-1.08}}}$
Ours $\textbf{28.45}^{\textbf{{\color[rgb]{0,0.88,0}+2.86}}}$ $\textbf{44.30}^{\textbf{{\color[rgb]{0,0.88,0}+0.05}}}$ $\textbf{27.29}^{\textbf{{\color[rgb]{0,0.88,0}+2.76}}}$ $\text{44.52}^{\text{{\color[rgb]{0.68,0,0}-0.53}}}$
@@
ALSO $\text{30.72}^{\text{{\color[rgb]{0,0.88,0}+0.91}}}$ $\text{47.34}^{\text{{\color[rgb]{0.68,0,0}-1.95}}}$ $\text{28.06}^{\text{{\color[rgb]{0.68,0,0}-1.75}}}$ $\text{48.10}^{\text{{\color[rgb]{0.68,0,0}-1.19}}}$
@@
Ours $\textbf{30.58}^{\textbf{{\color[rgb]{0,0.88,0}+0.78}}}$ $\textbf{49.80}^{\textbf{{\color[rgb]{0,0.88,0}+0.51}}}$ $\textbf{30.58}^{\textbf{{\color[rgb]{0,0.88,0}+0.78}}}$ $\textbf{49.80}^{\textbf{{\color[rgb]{0,0.88,0}+0.51}}}$
Table 1: Few-shot MOS results on the nuScenes dataset. The best performance in each column is in bold. The metric used for model selection is in bold in the table header.
5.3 Transferring Results on SemanticKITTI
To evaluate the transferability of the learned spatiotemporal representations, we fine-tune the pre-trained 4D UNet encoder on the SemanticKITTI dataset. We train the model on 10%/20%50% random samples from SemanticKITTI train set for enough epochs until convergence. As demonstrated in Tab. 2, TOP shows the best transferability on different LiDAR setups among all baselines. Fig. 5 displays the results on the SemanticKITTI dataset.
Data Pretrain Best $\textbf{mIoU}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
$\textbf{mIoU}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{mIoU}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
@@ 5.3 Transferring Results on SemanticKITTI
4Docc $-28.68\text{8.54}^{\text{{\color[rgb]{0.68,0,0}{-28.68}}}}$ $-27.87\text{13.35}^{\text{{\color[rgb]{0.68,0,0}{-27.87}}}}$ $-25.82\text{8.15}^{\text{{\color[rgb]{0.68,0,0}{-25.82}}}}$ $\text{13.45}^{\text{{\color[rgb]{0.68,0,0}-29.36}}}$
Uno $-0.57\text{36.66}^{\text{{\color[rgb]{0.68,0,0}{-0.57}}}}$ $-6.19\text{35.03}^{\text{{\color[rgb]{0.68,0,0}{-6.19}}}}$ $-6.15\text{27.83}^{\text{{\color[rgb]{0.68,0,0}{-6.15}}}}$ $\text{37.41}^{\text{{\color[rgb]{0.68,0,0}-5.41}}}$
Ours $\textbf{44.68}^{\textbf{{\color[rgb]{0,0.88,0}+7.46}}}$ $\textbf{43.29}^{\textbf{{\color[rgb]{0,0.88,0}+2.07}}}$ $\textbf{43.36}^{\textbf{{\color[rgb]{0,0.88,0}+9.39}}}$ $\textbf{49.37}^{\textbf{{\color[rgb]{0,0.88,0}+6.55}}}$
@@
4Docc $-21.80\text{25.61}^{\text{{\color[rgb]{0.68,0,0}{-21.80}}}}$ $-11.55\text{35.83}^{\text{{\color[rgb]{0.68,0,0}{-11.55}}}}$ $-20.16\text{25.23}^{\text{{\color[rgb]{0.68,0,0}{-20.16}}}}$ $\text{36.24}^{\text{{\color[rgb]{0.68,0,0}-12.53}}}$
@@
Ours $\textbf{58.45}^{\textbf{{\color[rgb]{0,0.88,0}+11.04}}}$ $\textbf{50.21}^{\textbf{{\color[rgb]{0,0.88,0}+2.82}}}$ $\textbf{58.45}^{\textbf{{\color[rgb]{0,0.88,0}+13.06}}}$ $\textbf{50.21}^{\textbf{{\color[rgb]{0,0.88,0}+1.44}}}$
@@
ALSO $\text{49.82}^{\text{{\color[rgb]{0.68,0,0}-3.32}}}$ $\text{57.07}^{\text{{\color[rgb]{0,0.88,0}+0.72}}}$ $\text{44.27}^{\text{{\color[rgb]{0.68,0,0}-3.93}}}$ $\text{60.25}^{\text{{\color[rgb]{0,0.88,0}+1.04}}}$
@@
Uno $\text{53.61}^{\text{{\color[rgb]{0,0.88,0}+0.48}}}$ $\text{54.97}^{\text{{\color[rgb]{0.68,0,0}-1.38}}}$ $\textbf{52.73}^{\textbf{{\color[rgb]{0,0.88,0}+4.53}}}$ $\text{57.06}^{\text{{\color[rgb]{0.68,0,0}-2.14}}}$
Ours $\textbf{55.81}^{\textbf{{\color[rgb]{0,0.88,0}+2.67}}}$ $\textbf{58.85}^{\textbf{{\color[rgb]{0,0.88,0}+2.50}}}$ $\text{52.11}^{\text{{\color[rgb]{0,0.88,0}+3.91}}}$ $\textbf{61.55}^{\textbf{{\color[rgb]{0,0.88,0}+2.35}}}$
Table 2: Cross-dataset transferring results on the SemanticKITTI MOS benchmark. The best performance in each column is in bold. The metric used for model selection is in bold in the table header.
@@ 5.4 Semantic Segmentation Results on nuScenes
We also assess how well our pre-training generalizes to a new downstream task by conducting a few-shot semantic segmentation experiment on the nuScenes dataset. We fine-tune the pre-trained encoders for a sufficient number of epochs using 10% of the train split sequences. As shown in Tab. 3, TOP shows a +0.42% mIoU improvement compared to the training-from-scratch baseline, while the performance of other methods notably declines. We suspect that the performance drop of 4Docc and UnO may be due to the complex “flow” information causing network overfitting, as discussed in Sec. 1. This experiment demonstrates the potential generalization capability of TOP pre-training across different downstream tasks.
Pretrain No ALSO 4Docc UnO Ours
mIoU 50.55 $50.26^{\textbf{{\color[rgb]{0.68,0,0}-0.29}}}$ $47.49^{\textbf{{\color[rgb]{0.68,0,0}-3.06}}}$ $48.06^{\textbf{{\color[rgb]{0.68,0,0}-2.49}}}$ $\textbf{50.97}^{\textbf{{\color[rgb]{0,0.88,0}+0.42}}}$
@@ 5.5 Ablation Study on Training Objectives
Objectives Best $\textbf{mIoU}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
$\textbf{mIoU}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{mIoU}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
$\mathcal{L}_{\text{overlap}}$ 24.559 34.785 22.639 35.350
@@
Table 4: Ablation study on two training objectives.
Figure 4: Visualization of the nuScenes MOS results.
Figure 5: Visualization of the SemanticKITTI MOS results.
@@ 6 Conclusion
In this paper, we propose a self-supervised pre-training method called Temporal Overlapping Prediction (TOP) for moving object segmentation (MOS). TOP predicts the occupancy states of temporal overlapping points to help the model learn spatiotemporal representations for MOS. We evaluate the effectiveness of TOP through few-shot experiments on nuScenes and cross-dataset transfer experiments on SemanticKITTI. TOP consistently improves both the object-level metric and the point-level metric, demonstrating strong transferability across LiDAR setups. Additionally, the improved performance in semantic segmentation suggests potential generalization to more downstream tasks. Further investigation of the effectiveness of TOP in other temporal tasks such as object tracking is worthwhile.

### v2

@@ 5.2 Few-shot Results on nuScenes
We evaluate few-shot MOS on nuScenes. We first pre-train an encoder on the entire training split. Then, the encoder is fine-tuned for point-wise MOS prediction using a binary Cross-Entropy loss on randomly sampled subsets (5%, 10%, 20%, 50%) of the nuScenes training sequences. All models are trained to convergence. For evaluation, we select the best-performing checkpoint for each of our two metrics separately. This model selection is performed on a validation set created by randomly partitioning the official nuScenes training split. The final results are reported on the official nuScenes validation split (used as our test set) in Tab. 1.
TOP outperforms the training-from-scratch baseline across all data subsets. It demonstrates substantial gains in $\text{Recall}_{\text{obj}}$, achieving up to 12.17% relative improvement (from 24.98% to 28.03%) on the 10% data subset. While the gains in $\text{IoU}_{\text{w/o}}$ are more modest, we attribute this to TOP’s enhanced ability to detect small moving objects, such as pedestrians and cyclists. These objects significantly contribute to recall but constitute a small fraction of the total points, thus having a lesser impact on IoU.
Data Pretrain Best $\textbf{Recall}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
$\textbf{Recall}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{Recall}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
@@
UnO $\text{25.23}^{\text{{\color[rgb]{0,0.6,0}+0.96}}}$ $\text{34.31}^{\text{{\color[rgb]{0.68,0,0}-0.65}}}$ $\text{22.62}^{\text{{\color[rgb]{0.68,0,0}-0.58}}}$ $\text{34.90}^{\text{{\color[rgb]{0.68,0,0}-0.15}}}$
TOP $\textbf{25.75}^{\textbf{{\color[rgb]{0,0.6,0}+1.48}}}$ $\textbf{35.94}^{\textbf{{\color[rgb]{0,0.6,0}+0.98}}}$ $\textbf{25.75}^{\textbf{{\color[rgb]{0,0.6,0}+2.56}}}$ $\textbf{35.94}^{\textbf{{\color[rgb]{0,0.6,0}+0.90}}}$
@@
ALSO $\text{27.32}^{\text{{\color[rgb]{0,0.6,0}+2.34}}}$ $\text{35.74}^{\text{{\color[rgb]{0.68,0,0}-0.70}}}$ $\text{22.47}^{\text{{\color[rgb]{0.68,0,0}-0.51}}}$ $\textbf{38.33}^{\textbf{{\color[rgb]{0,0.6,0}+0.27}}}$
@@
UnO $\text{26.38}^{\text{{\color[rgb]{0,0.6,0}+1.40}}}$ $\text{36.56}^{\text{{\color[rgb]{0,0.6,0}+0.12}}}$ $\text{22.00}^{\text{{\color[rgb]{0.68,0,0}-0.98}}}$ $\text{37.17}^{\text{{\color[rgb]{0.68,0,0}-0.89}}}$
TOP $\textbf{28.03}^{\textbf{{\color[rgb]{0,0.6,0}+3.04}}}$ $\textbf{36.95}^{\textbf{{\color[rgb]{0,0.6,0}+0.51}}}$ $\textbf{24.69}^{\textbf{{\color[rgb]{0,0.6,0}+1.72}}}$ $\text{38.06}^{\text{{\color[rgb]{0,0.6,0}+0.00}}}$
@@
ALSO $\text{26.68}^{\text{{\color[rgb]{0,0.6,0}+1.08}}}$ $\text{40.84}^{\text{{\color[rgb]{0.68,0,0}-3.41}}}$ $\text{25.81}^{\text{{\color[rgb]{0,0.6,0}+1.28}}}$ $\text{42.33}^{\text{{\color[rgb]{0.68,0,0}-2.72}}}$
4DOcc $\text{26.01}^{\text{{\color[rgb]{0,0.6,0}+0.42}}}$ $\text{42.21}^{\text{{\color[rgb]{0.68,0,0}-2.04}}}$ $\text{24.58}^{\text{{\color[rgb]{0,0.6,0}+0.06}}}$ $\text{43.11}^{\text{{\color[rgb]{0.68,0,0}-1.94}}}$
UnO $\text{28.14}^{\text{{\color[rgb]{0,0.6,0}+2.55}}}$ $\text{43.63}^{\text{{\color[rgb]{0.68,0,0}-0.62}}}$ $\text{26.78}^{\text{{\color[rgb]{0,0.6,0}+2.26}}}$ $\text{43.97}^{\text{{\color[rgb]{0.68,0,0}-1.08}}}$
TOP $\textbf{28.45}^{\textbf{{\color[rgb]{0,0.6,0}+2.86}}}$ $\textbf{44.30}^{\textbf{{\color[rgb]{0,0.6,0}+0.05}}}$ $\textbf{27.29}^{\textbf{{\color[rgb]{0,0.6,0}+2.76}}}$ $\text{44.52}^{\text{{\color[rgb]{0.68,0,0}-0.53}}}$
@@
ALSO $\text{30.72}^{\text{{\color[rgb]{0,0.6,0}+0.91}}}$ $\text{47.34}^{\text{{\color[rgb]{0.68,0,0}-1.95}}}$ $\text{28.06}^{\text{{\color[rgb]{0.68,0,0}-1.75}}}$ $\text{48.10}^{\text{{\color[rgb]{0.68,0,0}-1.19}}}$
@@
TOP $\textbf{30.58}^{\textbf{{\color[rgb]{0,0.6,0}+0.78}}}$ $\textbf{49.80}^{\textbf{{\color[rgb]{0,0.6,0}+0.51}}}$ $\textbf{30.58}^{\textbf{{\color[rgb]{0,0.6,0}+0.78}}}$ $\textbf{49.80}^{\textbf{{\color[rgb]{0,0.6,0}+0.51}}}$
Table 1: Few-shot MOS results on nuScenes. The best performance in each column is in bold. The metric used for model selection is in bold in the table header.
5.3 Few-shot Transfer Results on SemanticKITTI
We evaluate the transferability of our learned spatiotemporal representations by fine-tuning the pre-trained encoder on SemanticKITTI in a few-shot setting. Specifically, we use random subsets of 10%, 20%, and 50% of the official training set to fine-tune the model until convergence. We report the final results on the official SemanticKITTI validation set. As shown in Tab. 2, TOP demonstrates superior transferability, significantly outperforming the training-from-scratch baseline and other pre-training baselines across all data subsets. We present qualitative MOS results on nuScenes and SemanticKITTI in Fig. 4 and Fig. 5, respectively.
Data Pretrain Best $\textbf{Recall}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
$\textbf{Recall}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{Recall}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
@@ 5.3 Few-shot Transfer Results on SemanticKITTI
[…]
TOP $\textbf{44.68}^{\textbf{{\color[rgb]{0,0.6,0}+7.46}}}$ $\textbf{43.29}^{\textbf{{\color[rgb]{0,0.6,0}+2.07}}}$ $\textbf{43.36}^{\textbf{{\color[rgb]{0,0.6,0}+9.39}}}$ $\textbf{49.37}^{\textbf{{\color[rgb]{0,0.6,0}+6.55}}}$
@@
[…]
@@
TOP $\textbf{58.45}^{\textbf{{\color[rgb]{0,0.6,0}+11.04}}}$ $\textbf{50.21}^{\textbf{{\color[rgb]{0,0.6,0}+2.82}}}$ $\textbf{58.45}^{\textbf{{\color[rgb]{0,0.6,0}+13.06}}}$ $\textbf{50.21}^{\textbf{{\color[rgb]{0,0.6,0}+1.44}}}$
@@
ALSO $\text{49.82}^{\text{{\color[rgb]{0.68,0,0}-3.32}}}$ $\text{57.07}^{\text{{\color[rgb]{0,0.6,0}+0.72}}}$ $\text{44.27}^{\text{{\color[rgb]{0.68,0,0}-3.93}}}$ $\text{60.25}^{\text{{\color[rgb]{0,0.6,0}+1.04}}}$
@@
UnO $\text{53.61}^{\text{{\color[rgb]{0,0.6,0}+0.48}}}$ $\text{54.97}^{\text{{\color[rgb]{0.68,0,0}-1.38}}}$ $\textbf{52.73}^{\textbf{{\color[rgb]{0,0.6,0}+4.53}}}$ $\text{57.06}^{\text{{\color[rgb]{0.68,0,0}-2.14}}}$
TOP $\textbf{55.81}^{\textbf{{\color[rgb]{0,0.6,0}+2.67}}}$ $\textbf{58.85}^{\textbf{{\color[rgb]{0,0.6,0}+2.50}}}$ $\text{52.11}^{\text{{\color[rgb]{0,0.6,0}+3.91}}}$ $\textbf{61.55}^{\textbf{{\color[rgb]{0,0.6,0}+2.35}}}$
Table 2: Cross-dataset few-shot transfer results on SemanticKITTI. The best performance in each column is in bold. The metric used for model selection is in bold in the table header.
Figure 4: Qualitative results of the nuScenes MOS.
Figure 5: Qualitative results of the SemanticKITTI MOS.
@@ 5.4 Semantic Segmentation Results on nuScenes
We evaluate the generalization of our pre-training on a few-shot semantic segmentation task on nuScenes, by fine-tuning the pre-trained encoder using 10% of the training split. As shown in Tab. 3, TOP surpasses the training-from-scratch baseline by 0.42% mIoU at convergence. In contrast, models pre-trained with 4DOcc and UnO exhibit a notable performance drop. We attribute this drop to their overfitting to the complex “flow” information during their pre-training (Sec. 1). This result demonstrates the generalization potential of TOP pre-training across different downstream tasks.
Pretrain No ALSO 4DOcc UnO TOP
mIoU 50.55 $50.26^{\text{{\color[rgb]{0.68,0,0}-0.29}}}$ $47.49^{\text{{\color[rgb]{0.68,0,0}-3.06}}}$ $48.06^{\text{{\color[rgb]{0.68,0,0}-2.49}}}$ $\textbf{50.97}^{\textbf{{\color[rgb]{0,0.6,0}+0.42}}}$
@@ 5.5 Ablation Study on Training Objectives
Objectives Best $\textbf{Recall}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
$\textbf{Recall}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{Recall}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
$\mathcal{L}_{\text{overlap}}$ 24.56 34.79 22.64 35.35
@@
Table 4: Ablation study on training objectives.
@@ 6 Conclusion
In this paper, we propose Temporal Overlapping Prediction (TOP), a self-supervised pre-training method for MOS. By learning to predict the occupancy of temporal overlapping points, TOP develops powerful spatiotemporal representations for moving object segmentation. We evaluate its effectiveness through few-shot experiments on nuScenes and cross-dataset transfer experiments on SemanticKITTI. TOP consistently improves both object-level and point-level metrics, demonstrating strong transferability across LiDAR setups. Additionally, the improved performance in semantic segmentation suggests its potential to generalize to other downstream tasks. We believe investigating TOP for other temporal tasks, such as object tracking, is a promising avenue for future work.
Acknowledgement
The computations were performed using research computing facilities offered by Information Technology Services, the University of Hong Kong.
Supplementary Material
In this appendix, we describe a complementary formulation, the MOS labeling process, the cumulative distribution of moving objects and points, implementation details, the impact of ego-vehicle points on IoU, and potential negative social impacts. Finally, we present visual examples of the MOS results on the nuScenes and SemanticKITTI datasets in Fig. 7 and Fig. 8, respectively.
1 Formulation
As described in Sec. 3.2 Scenario-2 in the main paper, we provide the formula for the starting point $\mathbf{b}_{i}^{t_{0}}$ of the intersection segment. Since the spatial angle $\alpha_{i,j}$ is sufficiently small, we approximate that: (1) The two intersection segments are assumed to be equal in length, which is described as:
[…]
(2) The distance between the two starting points $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ is equal to the sum of beam radii at the two points, represented as follows:
[…]
We substitute Eq. 13 into Eq. 14 and solve for $\mathbf{b}_{i}^{t_{0}}||$:
[…]
2 MOS Labeling
As discussed in Sec. 5.1 of the main paper, the nuScenes object attributes do not precisely describe the motion state, for example, “cycle.with_rider” can be either moving or static. Furthermore, a proportion of objects lack attributes: $0.466\%$ of vehicles, $9.243\%$ of cycles, and $1.913\%$ of pedestrians. For our nuScenes MOS labeling, we calculate object speeds from their bounding box annotations. An object is classified as static if its speed is less than $\mu_{\text{sta}}$, and as moving if its speed exceeds $\mu_{\text{mov}}$. Objects with speeds between these thresholds are considered to have an unclear motion state and are thus classified as unknown. We use different speed thresholds for humans $(\mu_{\text{sta}}^{\text{hum}}=0.375,\mu_{\text{mov}}^{\text{hum}}=0.6)$, cycles $(\mu_{\text{sta}}^{\text{cyc}}=0.375,\mu_{\text{mov}}^{\text{cyc}}=1.0)$, and vehicles $(\mu_{\text{sta}}^{\text{veh}}=0.5,\mu_{\text{mov}}^{\text{veh}}=1.0)$.
3 Cumulative Distribution
Fig. 6 presents a statistical analysis of moving objects from the nuScenes train-val split (discussed in Sec. 4), plotting the cumulative distribution functions of moving object instances and the number of their scanned points. The distributions are analyzed with respect to object size, defined as the number of points per object (x-axis). The significant gap between the two curves highlights a strong imbalance: the vast majority of moving objects consist of very few points. This is evidenced by the data: while moving objects with 19 or fewer points comprise 75.22% of all moving objects, they contribute a mere 15.49% of the total moving points. This disparity grows, with the smallest 90.06% of objects collectively accounting for only 37.02% of the total points. Such a skewed distribution implies that a small number of point-rich objects can disproportionately dominate the conventional IoU metric.
Figure 6: Cumulative distribution of moving objects and points.
4 Implementation Details
method epochs batch size optimizer lr start lr max
TOP 50 8 AdamW $5.0\times 10^{-6}$ $5.0\times 10^{-5}$
ALSO 200 4 AdamW $\times 10^{-4} 1.0×10$$\times 10^{-3}$
4DOcc 50 2 AdamW $8.0\times 10^{-5}$ $8.0\times 10^{-4}$
UnO 50 16 AdamW $8.0\times 10^{-5}$ $8.0\times 10^{-4}$
Pre-training Settings. The detailed pre-training settings are shown in the table above. The learning rate (lr) follows a warmup-cosine schedule, increasing from “lr start” to “lr max” during the first 2% of iterations (the warmup stage), and then follows a cosine schedule for the remaining iterations. All pre-training baseline methods maintain their original settings, while 4DOcc uses a very small batch size due to the high GPU memory consumption caused by the 0.1 m dense grid map.
Fine-tuning Settings. The Adam optimizer is used for all the fine-tuning; the learning rate follows a step-decay schedule, starting from $1.0\times 10^{-4}$ and decreasing by a factor of $0.99$ after each iteration. For the nuScenes MOS experiments, the batch size is 4 for 5%, 10%, and 20% data subsets, while the batch size for the 50% data subset is 8. All methods are trained for over 400 epochs to ensure convergence. For SemanticKITTI cross-dataset transfer experiments, the batch size is 2 for all data subsets, and all methods are trained for over 300 epochs. For the nuScenes semantic segmentation experiment, all methods are trained for 300 epochs with a batch size of 4.
5 Impact of Ego-Vehicle Points on IoU Metric
As discussed in Sec. 4 of the main paper, the conventional IoU metric (denoted as IoU hereafter) includes ego-vehicle points, which inflates scores and masks actual perception performance on external moving objects. This issue is especially pronounced in the nuScenes dataset due to its high proportion of ego-vehicle points. To demonstrate this effect, we present nuScenes MOS results using IoU in Tab. 5. The results show that IoU scores are inflated by 1.5-2x compared to the ego-exclusive metric $\text{IoU}_{\text{w/o}}$. Furthermore, the apparent IoU performance gains are deceptive. These improvements are not from enhanced perception of the external environment, but from the far simpler task of ego-vehicle segmentation. Therefore, relying on this metric risks severely misrepresenting a model’s actual capability.
Data Pretrain Best $\textbf{Recall}_{\textbf{obj}}$
$\text{Recall}_{\text{obj}}$ $\text{IoU}_{\text{w/o}}$ IoU
10% No 24.98 36.44 68.89
TOP $\text{28.03}^{\text{{\color[rgb]{0,0.6,0}+3.04}}}$ $\text{36.95}^{\text{{\color[rgb]{0,0.6,0}+0.51}}}$ $\text{70.28}^{\text{{\color[rgb]{0,0.6,0}+1.39}}}$
20% No 25.59 44.25 63.68
TOP $\text{28.45}^{\text{{\color[rgb]{0,0.6,0}+2.86}}}$ $\text{44.30}^{\text{{\color[rgb]{0,0.6,0}+0.05}}}$ $\text{68.13}^{\text{{\color[rgb]{0,0.6,0}+4.45}}}$
Table 5: nuScenes MOS results using the conventional IoU metric.
6 Potential Negative Social Impacts
While the proposed self-supervised method for MOS reduces annotation costs for autonomous systems, potential societal impacts should be considered. The technology could unintentionally prioritize detection accuracy for dominant object classes (e.g., vehicles) over vulnerable road users like cyclists/pedestrians if trained on imbalanced datasets, potentially compromising safety in edge cases. The temporal nature of our approach might propagate motion prediction errors in complex urban scenarios, leading to hazardous decisions by autonomous vehicles. Furthermore, while addressing sensor bias through our new metric helps, residual geographic/cultural biases in training data (e.g., urban vs rural environments) could limit global applicability. The method could also be repurposed for surveillance systems that infringe on privacy. To mitigate these risks, we recommend: (1) Rigorous testing across diverse operational domains. (2) Implementing fairness-aware data sampling strategies. (3) Establishing ethical guidelines for secondary applications.
Figure 7: Qualitative results of the nuScenes MOS.
Figure 8: Qualitative results of the SemanticKITTI MOS.